# The contact process on dynamical random trees with degree dependence

### v1

@@ Abstract
The contact process is a simple model for the spread of an infection in a structured population. We investigate the case when the underlying structure evolves dynamically as a degree-dependent dynamical percolation model. Starting with a connected locally finite base graph we initially declare edges independently open with a probability that is allowed to depend on the degree of the adjacent vertices and closed otherwise. Edges are independently updated with a rate depending on the degrees and then are again declared open and closed with the same probabilities. We are interested in the contact process, where infections are only allowed to spread via open edges. Our aim is to analyze the impact of the update speed and the probability for edges to be open on the existence of a phase transition. For a general connected locally finite graph, our first result gives sufficient conditions for the critical value for survival to be strictly positive. Furthermore, in the setting of Bienaymé-Galton-Watson trees, we show that the process survives strongly with positive probability for any infection rate if the offspring distribution has a stretched exponential tail with an exponent depending on the percolation probability and the update speed. In particular, if the offspring distribution follows a power law and the connection probability is given by a product kernel and the update speed exhibits polynomial behaviour, we provide a complete characterisation of the phase transition.
@@ 1 Introduction
The long-term behaviour of the contact process is very well understood on fixed graphs such as the integer lattice and regular trees (see e.g. [14]). For finite graphs and also for infinite graphs with small infection rates $\lambda$ (possibly only $\lambda=0$) the infection will almost surely die out. (We will refer to this regime of extinction as the subcritical phase.) On infinite graphs, for larger values of $\lambda$ the infection may persist forever with positive probability. Within this regime there may still be a distinction between the possibility of weak survival referring to long-term persistence somewhere in the graph and strong survival referring to a positive probability for the infection to return infinitely often to its starting point.
Since fixed graphs are not particularly well suited for describing a complex interaction structure the contact process has more recently been considered on (finite and infinite) random graphs, in particular on random graphs with heavy-tailed degree distributions, where so-called hubs or small-world phenomena occur. Chatterjee and Durrett [5] were the first to highlight the important role of these hubs and they showed that the contact process does not exhibit a subcritical phase on finite graphs with a power-law degree distribution. In the setting of finite graphs this means that for all positive infection rates starting from all vertices infected, with high probability as the number of vertices tends to infinity, the process maintains a positive density of infected vertices for an exponentially long time (in the size of the graph). Of particular interest for our work are the results of Bhamidi et al. [2] and Durrett and Huang [9] who establish that a contact process on a (supercritical) Bienaymé-Galton-Watson (BGW) tree (on the event of the tree being of infinite size) exhibits a subcritical phase if and only if the offspring distribution has exponential moments. Recently, these results were extended by [4] to include an inhomogeneous model with random transmission rates.
@@
In this general setting, our first main result states that the model exhibits a subcritical phase if the penalisation is strong enough and the update mechanism fast enough. In other words, we derive sufficient conditions for the connection probabilities $p$ and the update speed $v$ under which the contact process goes extinct almost surely for a small enough infection rate. Furthermore, in this context, we show that the process does not explode in finite time almost surely for every infection rate. As an example, we show that these conditions are satisfied in certain parameter regimes for connection probabilities given by a product or maximum kernel.
The second main result involves a comparison with the penalised version of the contact process introduced recently in [1]. In particular, in the case when the underlying graph is a supercritical BGW tree, this comparison result together with results obtained in [1], enables us to deduce the absence of a subcritical phase in certain parameter regions. The third main result focuses on dynamical supercritical BGW trees and establishes sufficient conditions on the offspring distribution as well as on the $p$ and $v$ such that the contact process is supercritical on infinite trees for all positive infection rates. In particular, on dynamical supercritical BGW trees with a stretched exponential offspring distribution with a certain exponent (depending on $p$ and $v$), the contact process is always supercritical. Additionally, we show that immunisation does not occur in this setting in a large parameter region, meaning that survival is possible on infinite trees for infection rates that are large enough.
@@ 2 Model dynamics
for every $\{x,y\}\in E$. Throughout this paper, we always start the background process in its stationary distribution which means that for any fixed time $t\geq 0$ the probability that an edge is open ($\mathbb{P}(\{x,y\}\in\mathbf{B}_{t})=p(d_{x},d_{y})$ for all $\{x,y\}\in E$ independently of other edges.
[…]
where $\lambda>0$ denotes the infection rate and ||$\cdot$|| ⋅ | denotes the the cardinality of the set. On general graphs it is not a priori clear that the infection process is well defined. See Subsection 5.1 for a precise definition and a dicussion of this issue.
@@
$\kappa_{1}n^{-\alpha}$ $\leq p(n,m)\leq\kappa_{2}n^{-\alpha},$ (2.3)
$\nu_{1}n^{\eta}$ $\leq v(n,m)\leq\nu_{2}n^{\eta},$
for all $n\geq m$. Broadly speaking the condition on the connection probability $p$ can be interpreted in the following way. Let $x$ be a vertex with $d_{x}=N$. Then the probability to share an open edge with a neighbour at some time $t$ is approximately proportional to $N^{-\alpha}$. Since there are in total $N$ neighbours it follows if $N$ is large that
|$|\{y\in\mathcal{V}:\{x,y\}\in\mathbf{B}_{t}\}|\approx cN^{1-\alpha}.$ (2.4)
for some constant $c\geq 0$. We note that condition (2.3) on the connection probability $p$ is similar to the conditions used recently in [11, see equations (1) and (2)] for the case of finite graphs as well as by [1] on infinite graphs.
@@
These function are natural examples which satisfy (2.3). Here the parameter $\sigma$ indicates how much influence the smaller degree of the two neighbours has on the connection probability. For example $\sigma=0$ corresponds to the maximum kernel, i.e. $p_{\alpha,0}(d_{x},d_{y})=1\wedge\kappa(d_{x}\vee d_{y})^{-\alpha}$, in this case only the larger of both vertices has a contribution. If we choose $\sigma=1$ we get the product kernel, i.e. $p_{\alpha,1}(d_{x},d_{y})=1\wedge\kappa(d_{x}d_{y})^{-\alpha}$. Here, both vertices contribute fully relative to their degree. Another reason to consider these type of functions is that especially the maximum and product kernel have already been considered in a number of related articles, for example [13, 11, 1]. This enables us to compare our result with other established results. See Section 4 for more details.
Now, we define the critical values for the infection parameter $\lambda$. For a fixed vertex $\rho$, denote by $\mathbb{P}^{\{\rho\}}$ the law of $(\mathbf{C},\mathbf{B})$ started with $\rho$ infected. Given the graph $G$, we define the critical value between extinction and weak survival (or global survival) by
@@
and the critical value between weak and strong survival (or local survival) by
$\lambda_{2}(G):=\inf\big{\{}\lambda\geq 0:\ \liminf_{t\to\infty}\mathbb{P}^{\{\rho\}}\big{(}\rho\in\mathbf{C}_{t}\big{)}>0\big{\}}.$
Note that $\lambda_{1}(G)\leq\lambda_{2}(G)$ and that the definition of both of these critical values is independent of the specific choice of $\rho\in G$. The main contribution of this paper is to understand the phase transitions of the contact process in our setting. More precisely, our main results give sufficient conditions on the parameter of the model such that we have $\lambda_{1}(G)>0$ (and also for $\lambda_{2}(G)<\infty$) or $\lambda_{2}(G)=\lambda_{1}(G)=0$.
@@ 3 Main results
$v(d_{x},d_{y})\geq\underline{v}>0,$ (3.1)
where $\underline{v}$ denotes the minimum of the range of $v$, i.e.
$\underline{v}:=\min\{v(d_{x},d_{y}):x,y\in V\}.$
@@
Consider the CPDG $(\mathbf{C},\mathbf{B})$ on the graph $G$ started with any finite set $C\subset V$ of infected vertices. Assume that condition (3.1) holds. We also assume that there exists a function $W:\mathbb{N}\to[1,\infty)$ and a constant $0<K<\infty$ such that for all $x\in V$,
@@
$\sum_{\{x,y\}\in E}\frac{p(d_{x},d_{y})}{v(d_{x},d_{y})^{2}}<\frac{K}{\underline{v}^{2}}.$ (3.3)
Then we have $\lambda_{1}(G)>0$. Furthermore, it follows that the process does not explode, i.e.
@@
Consider the CPDG $(\mathbf{C},\mathbf{B})$ on the graph $G$ started with any finite set $C\subset V$ of infected vertices. Assume that $v$ satisfies (2.3) with $\eta\geq 0$ and that one of the following two conditions holds:
@@
Then, $\lambda_{1}(G)>0$ and the process does not explode for any infection rate $\lambda>0$.
@@
where $\lambda>0$ is the infection rate. Intuitively, this penalised contact process arises as the limit model as the update speed tends to infinity, where each potential infection event is immediately proceeded by an update event leading to a thinning of the infections with probability $p(d_{x},d_{y})$. In the second part of the result, we will consider a sequence of CPDG for which the minimum of the range of $v$ tends to infinity. In order to state the result we thus emphasize the dependence on $v$ of the critical values for extinction-survival and weak-strong survival by writing $\lambda_{1}(v,G)$ and $\lambda_{2}(v,G)$. Also, let $\lambda_{1}^{p}(G)$ and $\lambda_{2}^{p}(G)$ denote the critical values for extinction-survival and weak-strong survival, respectively, of the penalised contact process.
Proposition 3.3.
Consider the CPDG $(\mathbf{C},\mathbf{B})$ on the graph $G$ started with any finite set $C\subset V$ of infected vertices. If $\lambda_{1}^{p}(G)=0$ (resp. $\lambda_{2}^{p}(G)=0$) then we have $\lambda_{1}(v,G)=0$ (resp. $\lambda_{2}(v,G)=0$) for all $v$ that satisfy condition (3.1). Furthermore, we have for a sequence of CPDG $(\mathbf{C}^{n},\mathbf{B}^{n})_{n\in\mathbb{N}}$ with update speeds $(v^{n})_{n\in\mathbb{N}}$ satisfying (3.1) with $\lim_{n\rightarrow\infty}\underline{v}^{n}=\infty$ that
Our next main result focuses on the case when the underlying graph is given by a supercritical BGW tree. Let us denote by $\mathcal{T}\sim\text{BGW}(\zeta)$ the BGW tree rooted at $\rho$ with offspring distribution $\zeta$. Furthermore, we denote the corresponding vertex set by $\mathcal{V}$ and the edge set by $\mathcal{E}$, i.e. $\mathcal{T}=(\mathcal{V},\mathcal{E})$. Note that in this setting $\mathbb{P}$ denotes the probability measure for which the underlying random tree $\mathcal{T}$ and the CDPG $(\mathbf{C},\mathbf{B})$ are jointly defined. Sometimes we consider a fixed realization of $\mathcal{T}$, in these cases we denote the conditional probability by $\mathbb{P}_{\mathcal{T}}(\,\cdot\,):=\mathbb{P}(\ \cdot\ |\ \mathcal{T})$. The degree of a vertex $x\in\mathcal{V}$ is a random variable given by
$D_{x}=\begin{cases}\zeta_{\rho}&\text{ if }x=\rho\\ \zeta_{x}+1&\text{ otherwise }\end{cases}$ (3.5)
@@
$\lambda_{2}(\mathcal{T})=\lambda_{2}:=\inf\big{\{}\lambda>0:\liminf_{t\to\infty}\mathbb{P}^{\{\rho\}}(\rho\in\mathbf{C}_{t})>0\big{\}}$
@@
The first part of following theorem gives a sufficient criterion for the process to always survive strongly, so that there is no nontrivial phase transition. The second part gives us a sufficient condition under which the critical values are finite.
@@
Consider the CPDG $(\mathbf{C},\mathbf{B})$ on the BGW tree $\mathcal{T}$ satisfying (3.6). Assume that the connection probability $p$ and the update speed $v$ satisfy (2.3).
@@
    $\limsup_{N\to\infty}\frac{\log\mathbb{P}(\zeta=N)}{N^{1-\alpha-2(\eta\vee 0)}}=0,$ (3.7)
@@
    $\limsup_{N\to\infty}\frac{\log\mathbb{P}(\zeta=N)}{N^{1-\alpha}}=0,$ (3.8)
@@
Theorem 3.4$(i)$ yields a sufficient condition for the lack of a phase transition when the update speed is small, i.e. when $2\eta<1-\alpha$ with $10\leq\alpha<$ 1. In the situation of relatively fast updates, i.e. $v$ satisfying (2.3) with $\eta\geq 0$, and the connection probabilities $p$ being of the form as in (2.5), we will see that Proposition 3.3 gives us a sufficient condition when the subcritical phase is absent in our model by a comparison with a penalised contact process as introduced in (3.4).
@@
$\limsup_{N\to\infty}\frac{\log\mathbb{P}(\zeta=N)}{N^{1-2\alpha}}=0,$ (3.9)
@@
Both Theorem 3.4$(i)$ and Proposition 3.5 give criteria guaranteeing that there is no phase transition, but it depends on the value of $\eta$ which result applies for a larger range of $\alpha$. This indicates a transition from a regime where for slow update speeds our model is closer to the static case to a regime for relatively fast speed which is closer to the penalised contact process. We discuss this and other aspects more in depth in Section 4.

### v2

@@ Abstract
The contact process is a simple model for the spread of an infection in a structured population. We investigate the case when the underlying structure evolves dynamically as a degree-dependent dynamical percolation model. Starting with a connected locally finite base graph we initially declare edges independently open with a connection probability that is allowed to depend on the degree of the adjacent vertices and closed otherwise. Edges are independently updated with a rate depending on the degrees and then are again declared open and closed with the same probabilities. We are interested in the contact process, where infections are only allowed to spread via open edges. Our aim is to analyse the impact of the update speed and the connection probability on the existence of a phase transition. For a general connected locally finite graph, our first result gives sufficient conditions for the critical value for survival to be strictly positive. Furthermore, in the setting of Bienaymé-Galton-Watson trees, we show that the process survives strongly with positive probability for any infection rate if the offspring distribution has a stretched exponential tail with an exponent depending on the connection probability and the update speed. In particular, if the offspring distribution follows a power law and the connection probability is given by a product kernel and the update speed exhibits polynomial behaviour, we provide a complete characterisation of the phase transition.
@@ 1 Introduction
The long-term behaviour of the contact process is very well understood on fixed graphs such as the integer lattice and regular trees (see e.g. [14]). For finite graphs and also for infinite graphs with small infection rates $\lambda$ (possibly only $\lambda=0$) the infection will almost surely die out. On infinite graphs, for larger values of $\lambda$ the infection may persist forever with positive probability. Within this regime there may still be a distinction between the possibility of weak survival referring to long-term persistence somewhere in the graph and strong survival which means that the infection returns infinitely often to its starting point.
Since fixed graphs are not particularly well suited for describing a complex interaction structure the contact process has more recently been considered on (finite and infinite) random graphs, in particular on random graphs with heavy-tailed degree distributions, where so-called hubs or small-world phenomena occur. The realisation that these hubs are important dates back at least to Berger et al. [2]. Later on, Chatterjee and Durrett [6] analyse the occurrence of these hubs to showthat the contact process does not exhibit a subcritical phase on finite graphs with a power-law degree distribution. In the setting of finite graphs this means that for all positive infection rates starting from all vertices infected, with high probability as the number of vertices tends to infinity, the process maintains a positive density of infected vertices for an exponentially long time (in the size of the graph). Of particular interest for our work are the results of Bhamidi et al. [3] and Huang and Durrett [8] who establish that a contact process on a (supercritical) Bienaymé-Galton-Watson (BGW) tree (on the event of the tree being of infinite size) exhibits a subcritical phase if and only if the offspring distribution has exponential moments. Recently, these results were extended by [5] to include an inhomogeneous model with random transmission rates.
@@
In this general setting, our first main result (see Theorem 3.1) states that the model exhibits a subcritical phase if the penalisation is strong enough and the update mechanism fast enough. In other words, we derive sufficient conditions for the connection probabilities $p$ and the update speed $v$ under which the contact process goes extinct almost surely for a small enough infection rate. Furthermore, in this context, we show that the process does not explode in finite time almost surely for every infection rate. As an example, we show that these conditions are satisfied in certain parameter regimes for connection probabilities given by a product or maximum kernel.
The second main result (see Theorem 3.3) involves a comparison with the penalised version of the contact process introduced recently in [1]. In particular, in the case when the underlying graph is a supercritical BGW tree, this comparison result together with results obtained in [1], enables us to deduce the absence of a subcritical phase in certain parameter regions. The third main result (see Theorem 3.4) focuses on dynamical supercritical BGW trees and establishes sufficient conditions on the offspring distribution as well as on the $p$ and $v$ such that the contact process is supercritical on infinite trees for all positive infection rates. In particular, on dynamical supercritical BGW trees with a stretched exponential offspring distribution with a certain exponent (depending on $p$ and $v$), the contact process is always supercritical. Additionally, we show that immunisation does not occur in this setting in a large parameter region, meaning that survival is possible on infinite trees for infection rates that are large enough.
@@ 2 Model dynamics
for every $\{x,y\}\in E$. Throughout this paper, we always start the background process in its stationary distribution which means that for any fixed time $t\geq 0$ the probability that an edge is open $\mathbb{P}(\{x,y\}\in\mathbf{B}_{t})=p(d_{x},d_{y})$ for all $\{x,y\}\in E$ independently of other edges.
$C\mapsto C\cup\{x\}$ $\text{ with rate }\quad\lambda|\{y\in C:\{x,y\}\in B\}|,$ (2.2)
@@
where $\lambda>0$ denotes the infection rate and $|\cdot|$ denotes the cardinality of the set. On general graphs it is not a priori clear that the infection process is well defined. See Subsection 5.1 for a precise definition and a discussion of this issue.
@@
$\kappa_{1}n^{-\alpha}\leq p(n,m)$ $\leq\kappa_{2}n^{-\alpha}$ (2.3)
$\nu_{1}n^{\eta}\leq v(n,m)$ $\leq\nu_{2}n^{\eta},$ (2.4)
for all $n\geq m$. (Note that because of monotonicity $\kappa_{2}$ can in fact be chosen independently of $m$.) Broadly speaking the condition on the connection probability $p$ can be interpreted in the following way. Let $x$ be a vertex with $d_{x}=N$.Then the probability to share an open edge with a neighbour at some time $t$ is of lower or equal order than $N^{-\alpha}$. Since there are in total $N$ neighbours it follows
$|\{y\in\mathcal{V}:\{x,y\}\in\mathbf{B}_{t}\}|=O(N^{1-\alpha})$ (2.5)
as $N$ grows large. Note that for BGW trees, which we consider later, this quantity is with high probability exactly of order $N^{1-\alpha}$ as $N\to\infty$. We note that condition (2.3) on the connection probability $p$ is similar to the conditions used recently in [11, see equations (1) and (2)] for the case of finite graphs as well as by [1] on infinite graphs.
@@
These functions are natural examples which satisfy (2.3). Here the parameter $\sigma$ indicates how much influence the smaller degree of the two neighbours has on the connection probability. For example $\sigma=0$ corresponds to the maximum kernel, i.e. $p_{\alpha,0}(d_{x},d_{y})=1\wedge\kappa(d_{x}\vee d_{y})^{-\alpha}$, in this case only the larger of both vertices has a contribution. If we choose $\sigma=1$ we get the product kernel, i.e. $p_{\alpha,1}(d_{x},d_{y})=1\wedge\kappa(d_{x}d_{y})^{-\alpha}$. Here, both vertices contribute fully relative to their degree. Another reason to consider these types of functions is that especially the maximum and product kernel have already been considered in a number of related articles, for example [13, 11, 1]. This enables us to compare our result with other established results. See Section 4 for more details.
Now, we define the critical values for the infection parameter $\lambda$. For a fixed vertex $\rho$, denote by $\mathbb{P}^{\{\rho\}}$ the law of $(\mathbf{C},\mathbf{B})$ started with only $\rho$ infected. Given the graph $G$, we define the critical value between extinction and weak survival by
@@
and the critical value between weak and strong survival by
$\lambda_{2}(G):=\inf\big\{\lambda\geq 0:\ \mathbb{P}^{\{\rho\}}\big(\mbox{for any }s\geq 0\mbox{ there exists }t\geq s\,:\,\rho\in\mathbf{C}_{t}\big)>0\big\}.$
Note that $\lambda_{1}(G)\leq\lambda_{2}(G)$ and that the definition of both of these critical values is independent of the specific choice of $\rho\in G$. Thus, by the additivity of the contact process $\lambda_{1}(G)$ also has the same value if we start the contact process in any finite and nonempty set $C$ of infected sites. The main contribution of this paper is to understand the phase transitions of the contact process in our setting. More precisely, our main results give sufficient conditions on the parameter of the model such that we have $\lambda_{1}(G)>0$ (and also for $\lambda_{2}(G)<\infty$) or $\lambda_{2}(G)=\lambda_{1}(G)=0$.
@@ 3 Main results
$v(d_{x},d_{y})\geq\underline{v}:=\min\{v(d_{x},d_{y}):x,y\in V\}>0.$ (3.1)
@@
Consider the CPDG $(\mathbf{C},\mathbf{B})$ on the graph $G$ and assume that condition (3.1) holds. We also assume that there exists a function $W:\mathbb{N}\to[1,\infty)$ and a constant $0<K<\infty$ such that for all $x\in V$,
@@
$\sum_{\{x,y\}\in E}\frac{p(d_{x},d_{y})}{v(d_{x},d_{y})^{2}}\leq K.$ (3.3)
Then we have $\lambda_{1}(G)>0$. Furthermore, it follows that the process started with any finite set $C\subset V$ of infected vertices does not explode, i.e.
@@
Consider the CPDG $(\mathbf{C},\mathbf{B})$ on the graph $G$ and assume that $v$ satisfies (2.4) with $\eta\geq 0$ and that one of the following two conditions holds:
@@
Then, $\lambda_{1}(G)>0$ and the process started with any finite set $C\subset V$ of infected vertices does not explode for any infection rate $\lambda>0$.
@@
$C\mapsto C\cup\{x\}$ $\text{ with rate }\quad\lambda\sum_{y\in C,\{x,y\}\in E}p(d_{x},d_{y})$ (3.4)
@@
where $\lambda>0$ is the infection rate. Intuitively, this penalised contact process arises as the limit model as the update speed tends to infinity, where each potential infection event is immediately preceded by an update event leading to a thinning of the infections with probability $p(d_{x},d_{y})$. In the second part of the result, we will consider a sequence of CPDG for which the minimum of the range of $v$ tends to infinity. In order to state the result we thus emphasize the dependence on $v$ of the critical values for extinction-survival and weak-strong survival by writing $\lambda_{1}(v,G)$ and $\lambda_{2}(v,G)$. Also, let $\lambda_{1}^{p}(G)$ and $\lambda_{2}^{p}(G)$ denote the critical values for extinction-survival and weak-strong survival, respectively, of the penalised contact process.
Proposition 3.3.
Consider the CPDG $(\mathbf{C},\mathbf{B})$ on the graph $G$. If $\lambda_{1}^{p}(G)=0$ (resp. $\lambda_{2}^{p}(G)=0$) then we have $\lambda_{1}(v,G)=0$ (resp. $\lambda_{2}(v,G)=0$) for all $v$ that satisfy condition (3.1). Furthermore, we have for a sequence of CPDG $(\mathbf{C}^{n},\mathbf{B}^{n})_{n\in\mathbb{N}}$ with update speeds $(v^{n})_{n\in\mathbb{N}}$ satisfying (3.1) with $\lim_{n\rightarrow\infty}\underline{v}^{n}=\infty$ (and fixed connection probabilities $p$) that
Our next main result focuses on the case where the underlying graph is given by a supercritical BGW tree. Let us denote by $\mathcal{T}\sim\text{BGW}(\zeta)$ the BGW tree rooted at $\rho$ with offspring distribution $\zeta$. Furthermore, we denote the corresponding vertex set by $\mathcal{V}$ and the edge set by $\mathcal{E}$, i.e. $\mathcal{T}=(\mathcal{V},\mathcal{E})$. Note that in this setting $\mathbb{P}$ denotes the probability measure for which the underlying random tree $\mathcal{T}$ and the CDPG $(\mathbf{C},\mathbf{B})$ are jointly defined. Sometimes we consider a fixed realization of $\mathcal{T}$, in these cases we denote the conditional probability by $\mathbb{P}_{\mathcal{T}}(\,\cdot\,):=\mathbb{P}(\ \cdot\ |\ \mathcal{T})$. The degree of a vertex $x\in\mathcal{V}$ is a random variable given by
@@
$\lambda_{2}(\mathcal{T})$ $=\lambda_{2}:=\inf\big\{\lambda>0:\mathbb{P}^{\{\rho\}}(\forall s\geq 0\,\exists t\geq s\,:\,\rho\in\mathbf{C}_{t})>0\big\}$
@@
The first part of the following theorem gives a sufficient criterion for the process to always survive strongly, so that there is no nontrivial phase transition. The second part gives us a sufficient condition under which the critical values are finite.
@@
Consider the CPDG $(\mathbf{C},\mathbf{B})$ on the BGW tree $\mathcal{T}$ satisfying (3.6). Assume that the connection probability $p$ and the update speed $v$ satisfy (2.3) and (2.4).
@@
    $\limsup_{N\to\infty}\frac{\log\mathbb{P}(\zeta=N)}{N^{1-\alpha-2(\eta\vee 0)}}=0,\$(3.7)
@@
    $\limsup_{N\to\infty}\frac{\log\mathbb{P}(\zeta=N)}{N^{1-\alpha}}=0,\$(3.8)
@@
Theorem 3.4(i)(i) yields a sufficient condition for the lack of a phase transition when the system updates slowly, i.e. $\eta\leq 0$, or at moderate speed, i.e. $\eta\in(0,\tfrac{1}{2})$. In case the update speed is slow, we only need $\alpha<1$, but if it is moderate, a stricter assumption on $p$ is required, which is that $\alpha<(1-2\eta)$.
@@
$\limsup_{N\to\infty}\frac{\log\mathbb{P}(\zeta=N)}{N^{1-2\alpha}}=0,\$(3.9)
@@
If the update speed of the system is moderate or fast, where by fast we mean that $\eta\geq\tfrac{1}{2}$, and the connection probabilities $p$ is of the form as in (2.6), we will see that Proposition 3.5 gives us a sufficient condition when the subcritical phase is absent in our model by a comparison with a penalised contact process as introduced in (3.4). Here, the absence of a phase transition is shown if $\alpha<\tfrac{1}{2}$.
Both Theorem 3.4(i)(i) and Proposition 3.5 give criteria guaranteeing that there is no phase transition, but it depends on the value of $\eta$ which result applies for a larger range of $\alpha$. Indeed, which of the two provides the better result changes when the update speed of the system is moderate. This indicates that there might be a change of behaviour where our model is closer to the static case if the update speed is slow and for fast speed it behaves more like the penalised contact process. See Remark 3.6 for a more detailed discussion of this. Furthermore, we discuss this and other aspects in more depth in Section 4.